OpenStudy (anonymous):

sin 4u = 2 sin 2u cos 2u verify the identity

OpenStudy (anonymous):

@Mertsj @mebs

OpenStudy (anonymous):

@satellite73

OpenStudy (mertsj):

You have the identity sin 2x = 2sin x cos x

OpenStudy (mertsj):

Let x = 2u Then 2x = 4u Make the substitutions

OpenStudy (anonymous):

so it would be sin2x=2sinxcosx?

OpenStudy (anonymous):

and then from there what do i have to do just solve?

OpenStudy (anonymous):

so we have an identity \[\sin(A+B) = sinAcosB + cosAsinB\] we have sin(4u) = sin(2u + 2u) can you substitute and sovle?

OpenStudy (anonymous):

to solve would just leave me with cosx rigth ? @Mertsj @mebs

OpenStudy (anonymous):

no it wouldn't ... here it would look like this sin(2u + 2u) = sin2ucos2u + sin2ucos2u add them and you prove the identity.. its self explanatory.